Circular No. 8156

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 HT_{15} (LINEAR) An apparently asteroidal object reported by LINEAR (discovery obser-

vation below published on MPS 78496; prediscovery LINEAR observations published on MPS 80247; orbital elements on MPO 48372) has been found cometary by C. Hergenrother, who reports a diffuse coma of diameter 15′′ (and mag 18.6 within an aperture of radius 8′′) and a broad tail 60′′ long in p.a. 115^{o} on co-added 900-s R-band images taken on June 24.3 UT with the Mount Hopkins 1.2-m reflector.

2003 UT  α_{2000}  δ_{2000}  m_{2}
Apr. 26.26123  14^{h}34^{m}17.^{s}41  +21^{o}00′35′′.0  18.3

Epoch = 2003 May 1.0 TT

T = 2003 Apr. 17.7216 TT  ω = 124.^{o}0498
e = 0.419870  Ω = 81.4727 }2000.0
q = 2.671482 AU  i = 27.6701

a = 4.604970 AU n^{o} = 0.0997387 P = 9.88 years

2003 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 20 14^{h}05.^{m}65 +10^{o}43.^{′}1 2.116 2.706 114.^{o}9 19.^{o}9 18.0
30 14 07.11 + 8 05.1 2.229 2.717 107.7 20.9 18.1
July 10 14 10.65 + 5 25.4 2.353 2.731 100.7 21.5 18.2
20 14 16.04 + 2 47.0 2.484 2.745  93.8 21.7 18.4
30 14 23.07 + 0 11.9 2.620 2.762  87.1 21.5 18.5
Aug. 9 14 31.54 − 2 18.6 2.760 2.779  80.5 21.1 18.6
19 14 41.26 − 4 43.5 2.901 2.798  74.1 20.4 18.8
29 14 52.08 − 7 02.2 3.041 2.819  67.8 19.4 18.9

SUPERNOVA 2003gf IN MCG –04-52-26 Further to IAUC 8153, J. Graham and W. Li report the LOTOSS

discovery of an apparent supernova on unfiltered KAIT images taken on June 24.5 (mag ∼ 14.5) and 25.5 UT (mag ∼ 14.6). SN 2003gf is located at α = 22^{h}13^{m}40^{s}.94, δ = −21^{o}44′03′′.3 (equinox 2000.0), which is 19′′.2 east and 3′′.9 south of the diffuse nucleus of MCG –04-52-26. A KAIT image taken on 2002 Dec. 12.2 showed nothing at this position (limiting mag ∼ 19.0).

2003 June 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green